# Agent Design Pattern Catalogue: A Collection of Architectural Patterns for Foundation Model based Agents

### v3

@@ A Collection of Architectural Patterns
Foundation model-enabled generative artificial intelligence facilitates the development and implementation of agents, which can leverage distinguished reasoning and language processing capabilities to takes a proactive, autonomous role to pursue users’ goals. Nevertheless, there is a lack of systematic knowledge to guide practitioners in designing the agents considering challenges of goal-seeking (including generating instrumental goals and plans), such as hallucinations inherent in foundation models, explainability of reasoning process, complex accountability, etc. To address this issue, we have performed a systematic literature review to understand the state-of-the-art foundation model-based agents and the broader ecosystem. In this paper, we present a pattern catalogue consisting of 18 architectural patterns with analyses of the context, forces, and trade-offs as the outcomes from the previous literature review. The proposed catalogue can provide holistic guidance for the effective use of patterns, and support the architecture design of foundation model-based agents by facilitating goal-seeking and plan generation.
@@ 1 Introduction
In this regard, we present a catalogue of patterns for the design of foundation model-based agents in this paper. Please note that “agent” can be referred to i) AI acting on behalf of another entity, or; ii) AI that can take active roles or produces effect to achieve users’ goals. The former circumstance requires thorough analysis on governance perspective, while hereby, we claim that in this study, we focus on the second concept of “agents” that are capable of goal-seeking and plan generation. In software engineering, an architectural pattern is a reusable solution to a problem that occurs commonly within a given context in software design. Our pattern catalogue includes 18 patterns that were identified based on the study conducted by Lu et al. [5]. The intended audience of collected patterns is software architects and developers who are interested in FM-based agent design and implementation. The contributions of this paper include:
@@
The remainder of this paper is organised as follows. Section 2 introduces background knowledge and discusses related work. The methodology of this research is introduced in Section 3 and Section 4 presents each pattern in detail with our extended pattern template. Section 5 discusses the proposed pattern catalogue while Section 6 concludes the paper.
@@ 2 Background & Related Work
With the explosive growth of FMs, it is highly notable that FM-based agents come into the picture. AI agents are typically designed to operate a particular software environment. A single agent is able to take actions in a variety of three-dimensional virtual worlds [9]. Recently, there have been a lot of studies that present the architecture of their agents and broader AI systems [10, 11]. LangChain analysed the cognitive architecture of agents [12]. However, these architectures often focus only on certain components or schemes. For instance, Packer et al. [10] explicitly covered the memory design of the agent. Andrew Ng discussed reflection, tool use, planning and multi-agent collaboration [9] and provided a corresponding course [^8] https://www.deeplearning.ai/short-courses/ai-agentic-design-patterns-with-autogen/. Jain presented four design patterns for compound AI systems, including retrieval augmented generation, conversational AI, multi-agent communication, and co-pilot [13]. Yan et al. [14] shared their experience of building products with LLMs and the best practices they summarised during this process. Hassan et al. [15] demonstrated a high-level structure of FMware, consisting of Agentware and Promptware, and identified the challenges of software engineering for FMware. Gao et al. [16] proposed the roadmap for designing biomedical AI agents, and illustrated the required components.
@@ 3 Methodology
Based on the reported findings, we delved into the analysis of identified patterns for building integrating FM-based agents. Through the SLR, we have identified a series of architectural design challenges in the development and implementation of systems with the integration of agents. We then further conducted extensive review on this topic, which includes grey literature, and real-world applications (through scrutinising official websites and documents available), for identifying the known uses. Combining the findings of both SLR and the additional review, in this paper, we report our findings on 18 extracted patterns.
@@
Incremental model querying Access the foundation model at each step of the plan generation process to provide supplementary context, improve reasoning certainty and explainability
@@
Agent adapter Connect the agent and tools by learning new interfaces and converting incompatible interfaces into expected ones, ensuring interoperability and adaptability, and reduce development cost.
Agent evaluator Provide interface to connect the agent and external tools for task completion, ensuring the functional suitability, adaptability with improved flexibility.
@@ 4 Pattern Catalogue for Foundation Model-based Agents
In this section, we present a pattern catalogue for FM-based agents by adopting the extended pattern template in [19]. It includes the pattern name, a short summary, usage context, a problem statement, a discussion on the forces leading to the problem difficulty, the solution and its consequences, and several examples of real-world known uses of the pattern. Table 1 offers an overview of the collected patterns.
@@
A generated plan should be reviewed to ensure its accuracy, usability, completeness, etc. Self-refection, cross-reflection, and human reflection can help the agent to collect feedback from different reflective entities, and refine the plan and reasoning steps accordingly. Afterwards, the agent can assign tasks to other narrow AI-based or non-AI systems, invoke external tools, and employ a set of agents for goal achievement by tool/agent registry. In particular, agents can work on the same task and finalise the results with voting-based, role-based, or debate-based cooperation. For instance, agents can act as different roles such as coordinator and worker. Agent adapter keeps learning the interfaces of different tools, and convert them into FM-friendly environment. Multimodal guardrails can be applied to manage and control the inputs/outputs of foundation models. Please note that we omit the detailed architecture of agent-as-a-worker, and pattern application in several interactions for the simplicity of this diagram, for instance, the invocation of non-agent-AI/non-AI systems and external tools services can all apply tool/agent registry, and agent-as-a-worker can all have their respective datastore or knowledge base. Finally, we claim that developers can evaluate the performance of agents at both design-time and runtime via agent evaluator.
@@ 4.2 Proactive Goal Creator
Solution: Fig. 4 illustrates a simple graphical representation of proactive goal creator. In addition to the prompts received from dialogue interface, and relevant context retrieved from memory, the proactive goal creator can anticipate users’ goals by sending requirements to detectors, which will then capture and return the user’s surroundings for further analysis and comprehension to generate the goals, for instance, identifying the user’s gestures through cameras, recognising application UI layout via screenshots, etc. The proactive goal creator should notify users about context capturing and other relevant issues with a low false positive rate, to avoid unnecessary interruptions. In addition, the captured environment information can be stored in the agent’s memory (or knowledge base) to establish “world models” [22, 23] to improve its ability to comprehend the real world.
@@
  • Prompt/response optimiser.
  • Proactive goal creator can first handle users’ inputs and transfer the goals and relevant context information to prompt/response optimiser for prompt refinement.
@@ 4.4 Retrieval Augmented Generation (RAG)
Solution: Fig. 6 illustrates a high-level graphical representation of retrieval augmented generation. RAG is a technique for enhancing the accuracy and reliability of agents with facts retrieved from other sources (internal or online data). The knowledge gaps where the agents are lacking in memory are filled using the parameterized knowledge generated in vector databases. For instance, during plan generation, specific steps may require information that is not within the original agent memory. The agent can retrieve information from the parameterized knowledge and use for planning, while the augmented response (i.e. plan) will be return back to the user. The retrieval process requires zero pretraining or fine-tuning of the model served by the agent which preserves the data privacy of local data, reduces training and computation costs, and also provides up-to-date and more precise information required. The retrieved local data can be sent back to the agent via prompts (need to consider the context window size), whereas the agent is able to process the information and generate plans via in-context learning. Currently there is a cluster of RAG techniques focusing on various enhancement aspects, data sources and applications [29], for instance, federated RAG [30], graph RAG [31], etc.
@@ 4.5 One-Shot Model Querying
Solution: Fig. 7 illustrates interactions between user and agent within one-shot model query. After a user specifies goals and constraints in one prompt, the agent will query the incorporated foundation model to generate a corresponding response (e.g. plan for action). The foundation model does not require multiple interactions to comprehend the context and requirements. In this manner, the agent can devise a multi-step plan to achieve a broad goal, and provide a holistic explanation for this plan without delving into detailed reasoning steps.
@@ 4.6 Incremental Model Querying
Solution: Fig. 8 illustrates interactions between the users and agent within incremental model querying. The agent could engage in a step-by-step reasoning process to develop the plan for goal achievement with multiple queries to the foundation model. Meanwhile, human feedback can be provided at any time to both the reasoning process and generated plan, and adjustments can be made accordingly during model querying. Please note that incremental model querying can rely on a reusable template, which guides the process through context injection or an explicit workflow/plan repository and management system.
@@ 4.7 Single-Path Plan Generator
Solution: Fig. 9 illustrates a high-level graphical representation of single-path plan generator. After receiving and comprehending users’ goals, the single-path plan generator can coordinate the creation of intermediate steps for other agents or tools and prioritise the tasks, to progressively lead towards goal accomplishment. Specifically, each step in this process is designed to have only a single subsequent step, such as Chain-of-Thought (CoT) [39]. Self-consistency is employed to confirm with the foundation model several times and select the most consistent answer as the final decision [40]. Please note that the generated plan may have different granularity based on the given goal that complex plan may incorporate multiple workflows, processes, tasks and fine-grained steps.
@@ 4.8 Multi-Path Plan Generator
Solution: Fig. 10 illustrates a high-level graphical representation of multi-path plan generator. Based on single-path plan generator, multi-path plan generator can create multiple choices at each step towards the achievement of goals. Users’ preferences may influence the subsequent intermediate steps, leading to different eventual plans. The employment of involved agents and tools will be adjusted accordingly. Tree-of-Thoughts [42] exemplifies this design pattern.
@@ 4.9 Self-Reflection
Solution: Fig. 11 depicts a high-level graphical representation of self-reflection. In particular, reflection is an optimization process formalised to iteratively review and refine the agent-generated response. The user prompts specific goals to the agent, which then generates a plan to accomplish users’ requirements. Subsequently, the user can instruct the agent to reflect on the plan and the corresponding reasoning process. The agent will review the response to identify and pinpoint the errors, then generate a refined plan and adjust its reasoning process accordingly. The finalised plan will be carried out step by step. Self-consistency [43] exemplifies this pattern.
@@ 4.10 Cross-Reflection
Solution: Fig. 11 includes a high-level graphical representation of cross-reflection. If an agent cannot generate accurate results or precise planning steps via reflecting its outputs, users can prompt the agent to query another agent which is specialised in reflection. The latter agent can review and evaluate the outputs and relevant reasoning steps of the original agent, and provide refinement suggestions. This process can be iterative until the reflective agent confirms the plan. In addition, multiple agents can be queried for reflection to generate comprehensive responses.
@@ 4.11 Human Reflection
Solution: Fig. 11 presents a high-level graphical representation of human-reflection. When a user prompts his/her goals and specified constraints, the agent first creates a plan consisting of a series of intermediate steps. The constructed plan and its reasoning process can be presented to the user for review, or sent to other human experts to validate the feasibility and usefulness. The user or expert can provide comments or suggestions to indicate which steps can be updated or replaced. The plan will be iteratively assessed and improved until it is approved by the user/expert.
@@ 4.12 Voting-based Cooperation
Solution: Fig. 12 illustrates how agents can cooperate to finalise a decision via votes. Specifically, an agent can first generate a candidate response to the user’s prompts, then it holds a vote in which different reflective suggestions are presented as choices. Additional agents are requested to submit their votes to select the most appropriate feedback according to their capabilities and experiences. In this circumstance, agents communicate in a centralised manner that the original agent will act as a coordinator. The voting result will be formalised and sent back to the original agent, who can refine the response accordingly before answering the user.
@@ 4.13 Role-based Cooperation
Solution: Fig. 13 illustrates a high-level graphical representation of role-based cooperation, where agents coordinate in a hierarchical scheme. In particular, an agent-as-a-planner can generate a multi-step plan by decomposing user’s goal into a chain of tasks. Subsequently, the agent-as-an-assigner can orchestrate task assignment, i.e., some tasks can be completed by the assigner itself, while other tasks can be delegated to specific agent-as-a-worker based on their capabilities and expertise. In addition, if there is no available agent, agent-as-a-creator can be invoked to create a new agent with a specific role, by providing necessary resources, clear objectives and initial guidance to ensure a seamless transition of tasks and responsibilities. Please note that more elaborate roles can be defined and assigned to the agents.
@@
  • Scalability. Agents of new roles can be employed or created anytime to refine the task workflow and extend the capability of the whole system..

### v4

@@ A Collection of Architectural Patterns
Foundation model-enabled generative artificial intelligence facilitates the development and implementation of agents, which can leverage distinguished reasoning and language processing capabilities to takes a proactive, autonomous role to pursue users’ goals. Nevertheless, there is a lack of systematic knowledge to guide practitioners in designing the agents considering challenges of goal-seeking (including generating instrumental goals and plans), such as hallucinations inherent in foundation models, explainability of reasoning process, complex accountability, etc. To address this issue, we have performed a systematic literature review to understand the state-of-the-art foundation model-based agents and the broader ecosystem. In this paper, we present a pattern catalogue consisting of 18 architectural patterns with analyses of the context, forces, and trade-offs as the outcomes from the previous literature review. We propose a decision model for selecting the patterns. The proposed catalogue can provide holistic guidance for the effective use of patterns, and support the architecture design of foundation model-based agents by facilitating goal-seeking and plan generation.
@@ 1 Introduction
In this regard, we present a catalogue of patterns for foundation model-based agents in this paper, aiming to address the identified issues via providing a holistic guidance to the design and development of different types of agents, and specifying different collaboration schemes between these agents. For instance, the goal creator patterns can clarify users’ intentions and instructions to avoid underspecification. A series of patterns for reflection can help identify and mitigate the uncertainties in agent-generated plans, while the explainability of agent reasoning process is improved by requesting an agent to reflect on its generated plan. Accountability can be preserved when agents participate in a vote where their identities and operations are all logged. Please note that “agent” can be referred to i) AI acting on behalf of another entity, or; ii) AI that can take active roles or produces effect to achieve users’ goals. The former circumstance requires thorough analysis on governance perspective, while hereby, we claim that in this study, we focus on the second concept of “agents” that are capable of goal-seeking and plan generation. In software engineering, an architectural pattern is a reusable solution to a problem that occurs commonly within a given context in software design. Our pattern catalogue includes 18 patterns that were identified based on the study conducted by Lu et al. [5]. The intended audience of collected patterns is software architects and developers who are interested in FM-based agent design and implementation. The contributions of this paper include:
@@
  • A decision model that can help architects structure the included patterns and make rational design decisions on foundation model based agents. We also share the experiences on pattern application in different research projects.
The remainder of this paper is organised as follows. Section 2 introduces background knowledge and discusses related work. The methodology of this research is introduced in Section 3 and Section 4 presents each pattern in detail with our extended pattern template. Section 5 illustrates a decision model for pattern selection, and discusses the insights we obtained in this research project, while Section 6 concludes the paper.
@@ 2 Background & Related Work
With the explosive growth of FMs, it is highly notable that FM-based agents come into the picture. AI agents are typically designed to operate a particular software environment. A single agent is able to take actions in a variety of three-dimensional virtual worlds [9]. Recently, there have been a lot of studies that present the architecture of their agents and broader AI systems [10, 11]. LangChain analysed the cognitive architecture of agents [12]. However, these architectures often focus only on certain components or schemes. For instance, Packer et al. [10] explicitly covered the memory design of the agent. Andrew Ng discussed reflection, tool use, planning and multi-agent collaboration [9] and provided a corresponding course [^8] https://www.deeplearning.ai/short-courses/ai-agentic-design-patterns-with-autogen/. Jain presented four design patterns for compound AI systems, including retrieval augmented generation, conversational AI, multi-agent communication, and co-pilot [13]. Zhou et al. [14] proposed a four-tiered hierarchical artificial society model for the “cyber-physical-social” aspects of adaptive AI agents: i) an autonomous layer for agent memory, behavior and decision, ii) an evolutionary layer for agent learning and heterogeneity, iii) an interactive layer for agent collaboration, competition and topological structure, and iv) an emergent layer for the overall environment, feedback, and intervention. Yan et al. [15] shared their experience of building products with LLMs and the best practices they summarised during this process. Hassan et al. [16] demonstrated a high-level structure of FMware, consisting of Agentware and Promptware, and identified the challenges of software engineering for FMware. Gao et al. [17] proposed the roadmap for designing biomedical AI agents, and illustrated the required components.
@@ 3 Methodology
Based on the reported findings, we delved into the analysis of identified patterns for building integrating FM-based agents. Through the SLR, we have identified a series of architectural design challenges in the development and implementation of systems with the integration of agents. We then further conducted extensive review on this topic, which includes grey literature, and real-world applications (through scrutinising official websites and documents available), for identifying the known uses as the implementation of our included patterns. Combining the findings of both SLR and the additional review, in this paper, we report our findings on 18 extracted patterns.
@@
Incremental model querying Access the foundation model at each step of the plan generation process to provide supplementary context, improve reasoning certainty and explainability.
@@
Agent adapter Provide interface to connect the agent and external tools for task completion, ensuring interoperability and adaptability, and reduce development cost.
Agent evaluator Perform testing to assess the agent regarding diverse requirements and metrics, ensuring the functional suitability, adaptability with improved flexibility.
@@ 4 Pattern Catalogue for Foundation Model-based Agents
In this section, we present a pattern catalogue for FM-based agents by adopting the extended pattern template in [20]. It includes the pattern name, a short summary, usage context, a problem statement, a discussion on the forces leading to the problem difficulty, the solution and its consequences, and several examples of real-world known uses of the pattern. Please note that for each included pattern, we provide a simplified graphical representation that highlights only the essential components necessary to explain pattern application. Detailed interactions between all agent components have been omitted for clarity. Table 1 offers an overview of the collected patterns.
@@
A generated plan should be reviewed to ensure its accuracy, usability, completeness, etc. Self-refection, cross-reflection, and human reflection can help the agent to collect feedback from different reflective entities, and refine the plan and reasoning steps accordingly. Afterwards, the agent can assign tasks to other narrow AI-based or non-AI systems, invoke external tools, and employ a set of agents for goal achievement by tool/agent registry. In particular, agents can work on the same task and finalise the results with voting-based, role-based, or debate-based cooperation. For instance, agents can act as different roles such as coordinator and worker. Agent adapter keeps learning the interfaces of different tools, and convert them into FM-friendly environment. Multimodal guardrails can be applied to manage and control the inputs/outputs of foundation models. Meanwhile, the employed agents will conduct respective reasoning, planning and execution process, which may require external systems via retrieval augmented generation, tool/agent registry, and agent adapter either. Please note that we omit the detailed architecture of agent-as-a-worker, and pattern application in several interactions for the clarity of this diagram, for instance, each agent-as-a-worker has its passive/proactive goal creator, prompt/response optimiser, and single/multi-path plan generator, etc. Finally, we claim that developers can evaluate the performance of agents at both design-time and runtime via agent evaluator.
@@ 4.2 Proactive Goal Creator
Solution: Fig. 4 illustrates a simple graphical representation of proactive goal creator. In addition to the prompts received from dialogue interface, and relevant context retrieved from memory, the proactive goal creator can anticipate users’ goals by sending requirements to detectors, which will then capture and return the user’s surroundings with multimodal context information for further analysis and comprehension to generate the goals, for instance, identifying the user’s gestures through cameras, recognising application UI layout via screenshots, etc. Please note the proactive goal creator should notify users about context capturing and other relevant issues with a low false positive rate, to avoid unnecessary interruptions. In addition, the captured environment information can be stored in the agent’s memory (or knowledge base) to establish “world models” [23, 24] to continuously improve its ability to comprehend the real world.
@@
  • Prompt/response optimiser. Proactive goal creator can first handle users’ inputs and transfer the goals and relevant context information to prompt/response optimiser for prompt refinement.
  • Multimodal guardrails. Multimodal guardrails can help process the multimodal data captured by proactive goal creator.
@@ 4.4 Retrieval Augmented Generation (RAG)
Solution: Fig. 6 illustrates a high-level graphical representation of retrieval augmented generation. RAG is a technique for enhancing the accuracy and reliability of agents with facts retrieved from other sources (internal or online data). The knowledge gaps where the agents are lacking in memory are filled using the parameterized knowledge generated in vector databases. For instance, after plan generation, specific tasks may require information that is not within the original agent memory. The agent can hence retrieve information from the parameterised knowledge and use for task completion, while the augmented responses will be returned back to the user after optimisation. In particular, the implementation of RAG consists of the following steps: i) Determine the data sources. ii) Define the data structure for indexing raw data (i.e., text, image, video and audio) into embeddings or knowledge graphs. iii) Given a certain query, task executor encodes this query and searches the knowledge base (e.g., vector store) to retrieve the most relevant information. iv) Task executor processes the obtained data through reranking and filtering to produce a more informed and accurate response.
The retrieval process requires zero pretraining or fine-tuning of the model served by the agent which preserves the data privacy of local data, reduces training and computation costs, and also provides up-to-date and more precise information required. The retrieved local data can be sent back to the agent via prompts (need to consider the context window size), whereas the agent is able to process the information and generate plans via in-context learning. Currently there is a cluster of RAG techniques focusing on various enhancement aspects, data sources and applications [30], for instance, federated RAG [31], graph RAG [32], etc. Further, Retrieval Interleaved Generation can be considered a related technique of RAG where the agent can dynamically access external knowledge throughout the response generation phase.
@@ 4.5 One-Shot Model Querying
Solution: Fig. 7 illustrates interactions between user and agent within one-shot model query. In this scenario, the agent queries the incorporated foundation model to generate a corresponding plan based on user-specified goals and constraints. The foundation model is only queried for once in regard to the user’s requirements (e.g. limited budget), to comprehend the provided inputs. In this manner, the agent can devise a multi-step plan to achieve a broad goal, and provide a holistic explanation for this plan without delving into detailed reasoning steps. Please note that this pattern is applicable when other components query the integrated foundation model.
@@
  • Multimodal guardrails. Multimodal guardrails serve as an intermediate layer, managing the inputs and outputs of model querying.
@@ 4.6 Incremental Model Querying
Solution: Fig. 8 illustrates interactions between the plan generation component and integrated foundation model with incremental model querying. The agent could engage in a step-by-step reasoning process to develop the plan for goal achievement with multiple queries to the foundation model. Meanwhile, human feedback can be provided at any time to both the reasoning process and generated plan, and adjustments can be made accordingly during model querying. The number of queries can be predefined in agent configuration or determined in user prompts. Please note that incremental model querying can rely on a reusable template, which guides the process through context injection or an explicit workflow/plan repository and management system. This pattern is applicable when other components query the integrated foundation model.
@@
  • Multimodal guardrails. Multimodal guardrails serve as an intermediate layer, managing the inputs and outputs of model querying.
@@ 4.7 Single-Path Plan Generator
Solution: Fig. 9 illustrates a high-level graphical representation of single-path plan generator. After receiving and comprehending users’ goals, the single-path plan generator can coordinate the creation of a plan for other agents or tools and prioritise the tasks, to progressively lead towards goal accomplishment. Specifically, the plan generation process requires inference and reasoning that whether intermediate steps are actionable and optimal. Each step in this process is designed to have only a single subsequent step, forming a linear and direct plan, such as Chain-of-Thought (CoT) [40]. Self-consistency is employed to confirm with the foundation model several times and select the most consistent answer as the final decision [41]. Please note that the generated plan may have different granularity based on the given goal that complex plans may incorporate multiple workflows, processes, tasks and fine-grained steps.
@@ 4.8 Multi-Path Plan Generator
Solution: Fig. 10 illustrates a high-level graphical representation of multi-path plan generator. Based on single-path plan generator, multi-path plan generator can create multiple choices at each step towards the achievement of goals, which requires the underlying foundation model to tease out the eligible and actionable activities for each choice in the previous step. Specifically, users’ preferences may influence the subsequent intermediate steps, leading to different eventual plans. The employment of involved agents and tools will be adjusted accordingly. Tree-of-Thoughts [43] exemplifies this design pattern.
@@ 4.9 Self-Reflection
Solution: Fig. 11 depicts a high-level graphical representation of self-reflection. In particular, reflection is an optimisation process formalised to iteratively review and refine the reasoning process and generated contents of the agent. The user prompts specific goals to the agent, which then generates a plan to accomplish users’ requirements. Subsequently, the user can instruct the agent to reflect on the plan and the corresponding reasoning process. The agent will backtrack the inference process to verify whether certain intermediate results are incorrect and hence misleading all subsequent steps, then adjust and align its reasoning process to create a refined plan accordingly. Such reflection processes and results can be saved in the agent’s memory for continuous learning. The finalised plan will be carried out step by step. Self-consistency [44] exemplifies this pattern.
@@ 4.10 Cross-Reflection
Solution: Fig. 11 includes a high-level graphical representation of cross-reflection. If an agent cannot generate accurate results or precise planning steps via reflecting its outputs, users can prompt the agent to query another agent which is specialised in reflection. The latter agent can review and evaluate the logged outputs and relevant reasoning steps of the original agent, and provide refinement suggestions. This process can be iterative until the reflective agent confirms the plan. In addition, multiple agents can be queried for reflection to generate comprehensive responses.
@@ 4.11 Human Reflection
Solution: Fig. 11 presents a high-level graphical representation of human-reflection. When a user prompts his/her goals and specified constraints, the agent first creates a plan consisting of a series of intermediate steps. The constructed plan and its reasoning process logs can be presented to the user for review, or sent to other human experts to validate the feasibility and usefulness. The user or expert can provide comments or suggestions to indicate which steps can be updated or replaced. The plan will be iteratively assessed and improved until it is approved by the user/expert.
@@ 4.12 Voting-based Cooperation
Solution: Fig. 12 illustrates how agents can cooperate to finalise a decision via votes. Specifically, an agent can first generate a candidate response to the user’s prompts, then it holds a vote in which different reflective suggestions are presented as choices. Additional agents are requested to submit their votes to select the most appropriate feedback according to their capabilities and experiences. In this circumstance, agents communicate in a centralised manner that the original agent will act as a coordinator. The voting result will be formalised and sent back to the original agent, who can refine the response accordingly before answering the user. Please note that the voting process can be implemented through various methods, e.g., direct communications between the agent-as-a-coordinator and other agents, blockchain-based smart contracts, etc. Moreover, the identity management of all participating agents is significant to ensure the traceability and verifiability of votes.
@@ 4.13 Role-based Cooperation
Solution: Fig. 13 illustrates a high-level graphical representation of role-based cooperation, where agents coordinate in a hierarchical scheme. In particular, agents can be assigned certain roles and hence establishing a workflow via persona specification, task definition, tool employment, and process orchestration. For example, an agent-as-a-planner can generate a multi-step plan by decomposing user’s goal into a chain of tasks. Subsequently, the agent-as-an-assigner can orchestrate task assignment, i.e., some tasks can be completed by the assigner itself, while other tasks can be delegated to certain agent-as-a-worker based on their domain-specific capabilities and expertise. In addition, if there is no available agent, agent-as-a-creator can be invoked to create a new agent with a specific role, by providing necessary resources, clear objectives and initial guidance to ensure a seamless transition of tasks and responsibilities. Please note that more elaborate roles can be defined and assigned to the agents.
@@
  • Scalability. Agents of new roles can be employed or created anytime to refine the task workflow and extend the capability of the whole system.